OpenStudy (anonymous):

find a point slope equation of the line having the given slope and containing the given point. m=3/4, (6,2) what is the equation of the line? y-?=?(x-?)

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

the equation needs to be in point-slope form so y-y1=m(x-x1) you been given a point and slope so the answer is going to be y-2=3/4(x-6)